OpenStudy (anonymous):

Tickets for the school play cost $12 for adults and $8 for students. On opening nigh 120 tickets were sold and $1300 was collected. How many tickets were sold?

OpenStudy (zepp):

Let the number of students tickets be x and the number of adults tickets be y 120 tickets were sold > x+y=120 $1300 was collected > 8x+12y=1300

OpenStudy (zepp):

Now you have the system, solve for x, and y.

OpenStudy (anonymous):

\(x+y=120\) Isolate x \(x=120-y\) Substitute x in the second equation\[\begin{align}8(120-y)+12y&=1300\\960-8y+12y&=1300 \\960+4y&=1300\\ 4y&=1300-960\\ 4y&=340\\ y&=340\div4\\y=85 \end{align}\] So 85 adults tickets were sold, what about student? Substitute 85 in the first equation \(x+y=120\\x+85=120\\x=120-85\\x=35\) 35 Students tickets and 85 Adults tickets were sold.